OpenStudy (anonymous):

How do I find the remaining functions with the given information? (Will give medal & fan) csc x = sqrt(17), tan x = - 1/4 sin x = ? cos x = ? sec x =? cot x = ?

OpenStudy (anonymous):

Draw a triangle and fill out all of the information provided.

OpenStudy (anonymous):

Where does a triangle play into this?

OpenStudy (anonymous):

/:

OpenStudy (anonymous):

It gives you all of the sides you need.

OpenStudy (anonymous):

Ok so which sides go where? And is it a right triangle or other triangle?

OpenStudy (anonymous):

it actually gives you more information than you need

OpenStudy (anonymous):

Even if it gives me more info than I need can I still get the answers from it?

OpenStudy (anonymous):

This is the solution: csc x = sqrt(17),,,,,,,,,,,,,,,,,,,,,, csc x = 1/sin x 1/sin x = sqrt17 sin x = 1/ sqrt17 sin x = opp/ hyp so in this case your hyp has a value of sqrt(17) and your opposite is 1. Use the pythagorean triangle to figure out the adjacent side.

OpenStudy (anonymous):

Use this information to construct your triangle.

OpenStudy (anonymous):

So then sin x = 0.2425...?

OpenStudy (anonymous):

use exact answers, not decimals

OpenStudy (anonymous):

exact answers...?

OpenStudy (anonymous):

that's the answer that i got...

OpenStudy (anonymous):

@satellite73 ...?

OpenStudy (anonymous):

If I round it to 0.25 therefore making it 1/4 is that an "exact answer"?

OpenStudy (loser66):

you are good. 0.2425

OpenStudy (loser66):

to me, I do it simpler

OpenStudy (loser66):



OpenStudy (loser66):